# Magnetic hardness of hexagonal and orthorhombic Fe3C, Co3C, (Fe-Co)3C, and their alloys with boron, nitrogen, and transition metals: A first-principles study

### v2

@@ III.3.3 Magnetic moments and magnetic hardness
After discussing the formation energy and Curie temperature, we will show how alloying with transition metals affect the magnetic moments and magnetic hardness, see Figs. 7, 8, and 9 for hexa-Fe11X1C4, ortho-Fe11X1C4, and ortho-Co11X1C4 alloys, respectively. In the majority of instances, clear patterns emerge with regard to the dependencies. Trends in magnetic moments are similar to those calculated previously for alloys of ortho-Fe3C [33], CeFe12 [76], monoborides [64], and Fe [85, 86, 87]. MAE trends, on the other hand, do not resemble previous findings.
@@
Figures 7(d) and 7(f) show the magnetocrystalline anisotropy energies (MAE and DE32) for hexa-Fe11X1C4 alloys. From all considered compositions, the system with Pt has the highest MAE of 1.73 MJ m-3. This value corresponds to the highest magnetic hardness of 0.96, see Fig. 7(e). Our previous results for transition metal alloying in the Fe matrix, also pointed to Pt as a metal that stands out for its exceptionally positive effect on MAE [87]. Unfortunately, the high price of platinum, is not in favor of potential applications.
@@
In Sec. III.2 we discussed (Fe/Co)3C alloys in which substituting takes place at the 3 sites (Fe and Co). An additional degree of freedom in controlling the composition of Fe3C-type alloys is substituting at the 2 (carbon) site [21, 88, 53]. Two natural candidates for such alloying are boron and nitrogen – carbon’s neighbors in the periodic table. Although Fe3(B,C) borocarbides have been synthesized since at least the 1950s [21], they have not received wider attention. However, replacing carbon with boron in Fe3(B,C) alloys leads to significant changes in properties. Nicholson showed experimentally that in the alloys range from Fe3C to Fe3B, the magnetic moment grows by about 10% and the Curie temperature linearly increases from 481 K to the impressive 824 K [21]. A similar significant increase in Curie temperature during alloying at the -site was also observed for the transition from Fe5PB2 to Fe5SiB2, where an alloying element with one less atomic number was also used [89]. Medvedeva et al. have shown by supercell method that B alloying increases the magnetic moment of ortho-Fe3(B,C) borocarbides and stabilizes them [53]. Ande and Sluiter, also using supercell calculations, tested the effect of substitution of Al, Si, P, or S at the C-site in ortho-Fe3C on the stability of solid solutions [50]. Moreover, the hexa-Fe3N nitride has a negative formation energy, hence it stabilizes the hexagonal nitrocarbides [29]. The hexa-Fe3(C,N) powder consisting of 5% N was previously synthesized [29] and classified as magnetically soft.
@@
As we said, the most promising MAEs are observed for orthorhombic alloys with full Co content, see Fig. 10. To take a closer look at this range, in a one-dimensional Fig. 11 we again show the dependence of MAE on element concentration at the 2 position, along with a plot of spin magnetic moment. At the center of B-C-N range is the ortho-Co3C compound, known from previous discussion as a promising composition for hard magnetic applications. Although raising the nitrogen concentration leads to an increase in MAE, for reasons related to the instability of the ortho-Co3N phase, a better approach may be to alloy ortho-Co3C with boron, which should both stabilize the system and raise its Curie temperature. Experimental Curie temperatures of ortho-Co3C [23] and ortho-Co3B [16] are 650 K and 750 K, respectively. Moreover, the dependence of the magnetic moment on the concentration of 2 atoms is non-intuitive, see Fig. 11(b). While in the range from B to C it is qualitatively consistent with the ten percent decrease in magnetic moment measured for the isostructural Fe3(B,C) phase [21], in the range from C to N the magnetic moment shows a surprising minimum. Furthermore, the calculated MAE of Co3B of less than 0.2 MJ m-3 is much lower than the corresponding room-temperature experimental value of 0.65 MJ m-3 [16]. The discrepancy is due to, among other things, the dependence of MAE on temperature, as the calculations were performed at 0 K.
In summary, the alloying at the 2 position, as discussed in this section, allows for the control of MAE, magnetic moment, and Curie temperature. Furthermore, the aforementioned observation that MAE can depend on the total number of electrons in the system can prove useful in the designing of new compositions of magnetically hard materials.

### v3

@@ III.3.3 Magnetic moments and magnetic hardness
After discussing the formation energy and Curie temperature, we will show how alloying with transition metals affect the magnetic moments and magnetic hardness, see Figs. 8, 9, and 10 for hexa-Fe11X1C4, ortho-Fe11X1C4, and ortho-Co11X1C4 alloys, respectively. In the majority of instances, clear patterns emerge with regard to the dependencies. Trends in magnetic moments are similar to those calculated previously for alloys of ortho-Fe3C [37], CeFe12 [85], monoborides [68], and Fe [94, 95, 96]. MAE trends, on the other hand, do not resemble previous findings.
@@
Figures 8(d) and 8(f) show the magnetocrystalline anisotropy energies (MAE and DE32) for hexa-Fe11X1C4 alloys. From all considered compositions, the system with Pt has the highest MAE of 1.73 MJ m-3. This value corresponds to the highest magnetic hardness of 0.96, see Fig. 8(e). Our previous results for transition metal alloying in the Fe matrix, also pointed to Pt as a metal that stands out for its exceptionally positive effect on MAE [96]. Unfortunately, the high price of platinum, is not in favor of potential applications.
@@
In Sec. III.2 we discussed (Fe/Co)3C alloys in which substituting takes place at the 3 sites (Fe and Co). An additional degree of freedom in controlling the composition of Fe3C-type alloys is substituting at the 2 (carbon) site [21, 97, 57]. Two natural candidates for such alloying are boron and nitrogen – carbon’s neighbors in the periodic table. Although Fe3(B,C) borocarbides have been synthesized since at least the 1950s [21], they have not received wider attention. However, replacing carbon with boron in Fe3(B,C) alloys leads to significant changes in properties. Nicholson showed experimentally that in the alloys range from Fe3C to Fe3B, the magnetic moment grows by about 10% and the Curie temperature linearly increases from 481 K to the impressive 824 K [21]. A similar significant increase in Curie temperature during alloying at the -site was also observed for the transition from Fe5PB2 to Fe5SiB2, where an alloying element with one less atomic number was also used [98]. Medvedeva et al. have shown by supercell method that B alloying increases the magnetic moment of ortho-Fe3(B,C) borocarbides and stabilizes them [57]. Ande and Sluiter, also using supercell calculations, tested the effect of substitution of Al, Si, P, or S at the C-site in ortho-Fe3C on the stability of solid solutions [54]. Moreover, the hexa-Fe3N nitride has a negative formation energy, hence it stabilizes the hexagonal nitrocarbides [33]. The hexa-Fe3(C,N) powder consisting of 5% C was previously synthesized [33] and classified as magnetically soft.
@@
As we said, the most promising MAEs are observed for orthorhombic alloys with full Co content, see Fig. 11. To take a closer look at this range, in a one-dimensional Fig. 12 we again show the dependence of MAE on element concentration at the 2 position, along with a plot of spin magnetic moment. At the center of B-C-N range is the ortho-Co3C compound, known from previous discussion as a promising composition for hard magnetic applications. Although raising the nitrogen concentration leads to an increase in MAE, for reasons related to the instability of the ortho-Co3N phase, a better approach may be to alloy ortho-Co3C with boron, which should both stabilize the system and raise its Curie temperature. Moreover, the dependence of the magnetic moment on the concentration of 2 atoms is non-intuitive, see Fig. 12(b). While in the range from B to C it is qualitatively consistent with the ten percent decrease in magnetic moment measured for the isostructural Fe3(B,C) phase [21], in the range from C to N the spin magnetic moment shows a surprising minimum. It comes from the minima of spin magnetic moments at both non-equivalent Co positions. Light is shed on the minima in magnetic moments by density analysis, which in the case of the FPLO code involves determining the number of electrons in local density functions. The excess charge on Co atom relative to the neutral cobalt atom is shown in Fig. 12(c). The electrons from the cobalt atoms transfer to the 2 atoms. The observed minimum in the magnetic moment correlates with the charge minimum. Additional calculations without spin polarization (non-magnetic) show that the charge minimum is primary to the magnetic moment. The theoretical minimum discussed here for the orthorhombic phase may not be experimentally verifiable, since experiments to date for Co3N1+δ [99] and (Fe,Co)3N [100] indicate that Co3N prefers the hexagonal structure.
Furthermore, the calculated MAE of Co3B of less than 0.2 MJ m-3 is much lower than the corresponding room-temperature experimental value of 0.65 MJ m-3 [16]. The discrepancy is due to, among other things, the dependence of MAE on temperature, as the calculations were performed at 0 K.
As we discussed earlier, the experimental Curie temperature values for ortho-Co3C were determined for nanoparticles and range from 498 K to 650 K [24, 25, 26, 27] and are also significantly higher than the value for solid ortho-Fe3C of 481 K [21]. Mean-field theory calculations, which customarily overestimate the Curie temperature, give for ortho-Co3C a value of 458 K, suggesting that we should actually expect an experimental about 300 K for ortho-Co3C. The experimental Curie temperatures of ortho-Co3B are between 710 K [76] and 750 K [16], much higher than for the isostructural carbide. In Fig. 13, we show the results of Curie temperature calculations in the ortho-Co3(B-C-N) concentration range. In agreement with the experiment, we observe a decrease in Curie temperature with the transition from Co3B to Co3C. However, for the transition from Co3C to Co3N, we observe a maximum in the Curie temperature. A similar maximum in was observed experimentally for -Fe-N system with increasing nitrogen concentration [101]. As we mentioned above, since Co3N seems to prefer the hexagonal phase, the presented results for the ortho-Co3(C-N) system will most probably remain only interesting theoretical considerations.
In summary, the alloying at the 2 position, as discussed in this section, allows for the control of MAE, magnetic moment, and Curie temperature. Alloying the considered carbides with boron leads to an increase in the Curie temperature by several tens of percent. The increase in Curie temperature during alloying with boron correlates with an increase in magnetic moment up to about 10%. For the more stable orthorhombic phases, the highest magnetic hardnesses () are found near the Co3C composition, while remaining similarly high until about 30% boron alloying. Furthermore, the observation that MAE can depend on the total number of electrons in the system can prove useful in the designing of new compositions of magnetically hard materials.